OpenStudy (anonymous):

Free Fall Question. In Bosnia ,the ultimate test of a young man's courage used to be to jump off a 400-year-old bridge into the River Neretva,23m below the bridge.(a)How long did the jump last? (b) How fast was the jumper traveling upon impact with the river?(c)If the speed of sound in air is 340m/s, how long after the jumper took off did a spectator on the bridge hear the splash?

OpenStudy (anonymous):

a) if there's no air drag, time = 2.1

OpenStudy (anonymous):

b) is it means velocity?

OpenStudy (anonymous):

what is the formula to get that answer?

OpenStudy (anonymous):

(b) is velocity

OpenStudy (anonymous):

ok. a) is a physics problem. maybe situation is that at first man maybe at rest so initial velocity is zero and in the free-fall equation which is h=vt-(gt^2)/2, where v is initial velocity, v = 0. so t = root(2h/g)

OpenStudy (anonymous):

so plug in g = 10 (approximate value)

OpenStudy (anonymous):

h=23 then you can get answer.

OpenStudy (anonymous):

maybe proof of the free-fall equation is hard to write down in this web. so i recommand you to flipping the physics book. it is at firtst part of the book

OpenStudy (anonymous):

2.144right just making sure I'm getting this :D

OpenStudy (anonymous):

in b), also can solve it by free fall equation

OpenStudy (anonymous):

in the free fall equation v' = v - gt (v=initial velocity v'=final velocity) so v=0 and you've gotten the time, so just plug in value

OpenStudy (anonymous):

2.144 is how long jump last answer to part a, I need to learn how to set problem up correctly and what to plug in I thank you so much physics is killing me :(

OpenStudy (anonymous):

and c) is solving by v=s/t , so at this question v=340m/s and s is 23m so you can get time easily

OpenStudy (anonymous):

if you study physics book just 1 hour you can get sense how to set up free fall problem in all situation